### v1

@@ 4.6 Matrix Extension for Multivariate Analysis
In applications, we often augment $\text{CM}(y,Z)$ with selected columns from $\text{CM}(Z,Z)$ to form an extended association profile. For a vertex $v$, this profile includes both $\text{cm}(y,z_{j})(v)$ for all features $j$ and $\text{cm}(z_{j},z_{k})(v)$ for pairs $(j,k)$ of interest. These profiles embed vertices into an association space where proximity reflects similarity in how outcome and features relate locally. Vertices with similar profiles form co-monotonicity cells, which we will explore in subsequent sections as a basis for geometric multiple testing and biclustering.
4.7 Computational Considerations and Implementation
The algorithmic complexity of computing co-monotonicity coefficients is linear in the number of edges for sparse graphs. For a single pair of functions $y$ and $z$ on a graph with $n$ vertices and $|E|$ edges, the algorithm iterates over each vertex, examines its incident edges, and accumulates weighted products. With bounded vertex degree (typical for k-nearest neighbor graphs with fixed $k$), this yields $O(nk)=O(|E|)$ time complexity. For $m$ features in matrix form, optimized computation requires $O(m\cdot|E|)$ time after initial graph preprocessing. Space complexity is $O(n+|E|)$ for storing the graph plus $O(nm)$ for the result matrix, which is optimal since the output has this size.
For large-scale applications with thousands of vertices and hundreds of features, we employ OpenMP parallelization while maintaining CRAN compliance. The matrix computation naturally parallelizes over features: each column of $\text{CM}(y,Z)$ can be computed independently once the $y$-dependent quantities are precomputed. We distribute columns across threads, with each thread maintaining its own accumulation buffers to avoid race conditions. The Makevars configuration detects OpenMP support at compile time and enables parallelization on systems where it is available, gracefully degrading to serial computation otherwise.
Special numerical considerations arise in edge cases. When the denominator $\sum_{u\in N(v)}w_{e}|\Delta_{e}y\cdot\Delta_{e}z|$ vanishes, all edge differences have opposite-signed products that exactly cancel, or all products are zero. We define the coefficient as zero in this case, reflecting absence of systematic directional relationship. For derivative weighting with $w_{e}=1/\Delta_{e}^{2}$, we guard against division by zero for edges with negligible length $\Delta_{e}<10^{-10}$, setting $w_{e}=0$ for such edges to exclude them from the computation. This arises rarely in practice but prevents numerical instability.
4.8 Applications and Interpretation
Co-monotonicity coefficients support diverse analytical tasks in high-dimensional data analysis. We outline several key applications that motivate the subsequent development of inference frameworks.
[…]
For feature selection in regression or classification problems, we compute $\text{CM}(y,Z)$ to identify features most concordant with outcomes. Unlike Pearson correlation which measures linear association globally, co-monotonicity assesses directional concordance locally while respecting graph geometry. Features with high co-monotonicity in particular regions may warrant investigation even if their global correlation is modest, as they may drive outcomes in specific contexts.
Within gradient flow cells established through validated basin computation, co-monotonicity provides natural association measures. Since the outcome $\hat{y}$ varies monotonically within each cell by construction, features that exhibit monotone co-variation with $\hat{y}$ within the cell are prime candidates for mechanistic relationships. We restrict co-monotonicity computation to edges within a cell, yielding cell-specific association coefficients that respect the partition structure.
For multivariate feature analysis, the tensor $\text{CM}(Z,Z)$ reveals how feature relationships vary spatially. In microbiome studies, bacterial taxa may exhibit synergistic co-occurrence in some samples and competitive exclusion in others. The vertex-wise feature co-monotonicity matrix $\text{CM}(Z,Z)[v,:,:]$ at each vertex $v$ captures this local correlation structure, enabling identification of context-dependent feature modules through clustering or network analysis.
Finally, co-monotonicity matrices provide embeddings for biclustering and geometric multiple testing. By treating rows of $\text{CM}(y,Z)$ or selected components of $\text{CM}(Z,Z)$ as feature vectors, we embed vertices into an association space where distances reflect similarity in how features relate to outcomes. Clustering this embedding induces a partition into co-monotonicity cells, which we develop formally in subsequent sections as an alternative to gradient flow cells for discovering structure.
These applications share a common theme: moving from global, homogeneous association measures to spatially-resolved, geometry-aware coefficients that respect the structure latent in high-dimensional data. The vertex-level formulation enables detection of regional patterns that global methods would average away, while the matrix extensions provide computational efficiency for high-dimensional feature spaces. We now proceed to develop the statistical inference framework necessary to distinguish genuine associations from sampling noise in these geometry-respecting measures.
@@ 6.7 Relationship Between Gradient Flow and Co-Monotonicity Cells
6.8 Summary
The geometric multiple testing framework via co-monotonicity cells addresses the fundamental challenge of discovering spatially heterogeneous multivariate associations. By embedding vertices into association space through co-monotonicity profiles and applying community detection to the induced graph, we partition both the sample space and feature space simultaneously, identifying regions where specific feature modules exhibit coherent relationships with outcomes. The Bayesian uncertainty quantification through posterior sampling provides credible regions and modules with formal probability interpretations, while the region-module coherence measure enables statistical inference that borrows strength across vertices and features within cells. This approach dramatically reduces the multiple testing burden from $n\times m$ vertex-feature pairs to $K\times M$ region-module coherences where $K\ll n$ and $M\ll m$, while revealing context-dependent association patterns that global methods would obscure and individual tests would lack power to detect. The comparison between gradient flow cells (outcome-driven partitioning) and co-monotonicity cells (association-driven partitioning) provides complementary perspectives on the geometric organization of high-dimensional data, with concordance between partitions suggesting mechanistic coherence and discordance revealing multiple pathways to similar outcomes.
@@ 7. Discussion
We have developed a geometric framework for statistical inference on high-dimensional structured data that addresses a fundamental challenge: when features associate with outcomes differently across regions of the sample space, global methods dilute these signals while local methods lack the power and interpretability to discover the underlying structure. Our approach transforms this challenge into a geometric problem, representing data through Riemannian graphs and discovering structure through two complementary decomposition strategies.
7.1 Summary of Contributions
The first strategy employs gradient flow analysis to partition the sample space based on outcome geometry. Given a smoothed estimate of conditional expectation on graph vertices, we compute discrete gradient trajectories through validated gradient flow that addresses the long edge problem via path monotonicity. Long edges connecting geometrically distant points often create artifacts in naive gradient flow computation, with trajectories jumping between basins. Our path monotonicity criterion validates edges by verifying that alternative paths exhibit consistent monotonic behavior, ensuring gradient flow follows the intrinsic data geometry rather than graph construction artifacts. The resulting basins of attraction partition vertices into gradient flow cells where outcomes vary monotonically, providing natural regions for regional statistical modeling.
The second strategy discovers structure through association patterns rather than outcome values. Co-monotonicity coefficients quantify directional concordance between functions at individual graph vertices, measuring whether two functions tend to change together or oppositely across incident edges. These vertex-level measures respect graph geometry through derivative weighting that normalizes by edge length, connecting to angular correlation of gradients in the continuous limit. By computing co-monotonicity between outcomes and features, as well as among features themselves, we construct association profiles that embed vertices into a high-dimensional space where geometric proximity reflects similarity in how features relate to outcomes. Community detection on graphs constructed in this association space induces biclustering, simultaneously partitioning vertices into regions and features into modules. The resulting co-monotonicity cells identify where specific feature combinations drive outcomes, revealing multivariate association structure that univariate methods cannot capture.
For statistical inference, we develop both classical and Bayesian approaches. Vertex-wise permutation testing provides distribution-free inference with exact finite-sample control of type I error rates, employing FDR correction and spatial thresholding to handle multiple comparisons. More fundamentally, we develop Bayesian inference through posterior sampling on weighted Laplacians, using Dirichlet resampling to perturb vertex masses and generate ensembles of smoothed estimates. This propagates uncertainty from estimation through association quantification, yielding credible intervals and posterior probabilities for effect sizes. The Bayesian framework extends naturally to biclustering: computing co-clustering probabilities across posterior samples quantifies uncertainty in the discovered partition structure itself, enabling formal inference on region-module coherences with strength borrowing across vertices and features within cells.
7.2 Implementation and Validation
The methods are implemented in the R package gflow, which provides computationally efficient C++ implementations of core algorithms with CRAN-compliant interfaces. Detailed implementation documentation, computational complexity analysis, performance optimization strategies, and comprehensive empirical validation on real datasets will be provided in subsequent versions of this manuscript. The package is designed for production use in high-dimensional biomarker studies, with particular emphasis on microbiome-outcome association analysis where the methods were originally developed.
7.3 Relationship to Existing Methodologies
@@
Our work builds on and extends several methodological traditions. Morse-Smale regression, introduced by Gerber et al., pioneered the application of discrete Morse theory to statistical regression on graphs. Our contributions advance this framework through robust conditional expectation estimation on density-aware Riemannian graphs (detailed in the companion paper on adaptive geometric regression), systematic treatment of spurious extrema through prominence filtering and basin overlap clustering, path monotonicity validation that resolves long edge artifacts through geometric rather than heuristic criteria, and co-monotonicity measures that provide alternatives to linear models within cells while enabling discovery of association-driven structure complementary to outcome-driven gradient flow decomposition.
@@
The framework situates within geometric data analysis rather than topological data analysis. While TDA emphasizes topological invariants like homology groups and persistence diagrams that remain unchanged under continuous deformations, our approach focuses on geometric properties like distances, angles, and geodesics that depend on specific metric structure. The Riemannian graphs with vertex and edge masses capture local density and geometry that spectral methods exploit for inference. The gradient flow respects this geometry through path monotonicity validation and through the Riemannian metric implicit in the normalized Laplacian. This geometric emphasis connects our work to manifold learning and spectral graph theory more than to persistent homology, though we acknowledge that the basin complex structure has topological aspects.
From a statistical perspective, our geometric decomposition provides data-driven stratification for inference. Stratified sampling and testing are well-established in statistics, typically requiring investigators to specify strata based on known covariates. Our contribution makes stratification itself a geometric inference problem: strata emerge from graph structure through gradient flow or co-monotonicity analysis rather than being predetermined. This connects to recent work on conditional independence testing and context-specific associations, but grounds these ideas in explicit geometric partitions of the sample space derived from the Riemannian structure.
Compared to standard correlation-based methods, co-monotonicity offers several advantages. Classical correlation measures linear association globally, treating all observations as equally related and summarizing with a single coefficient. Co-monotonicity provides spatially-resolved assessment of directional concordance, detecting where associations are strong or weak across the graph while accounting for edge geometry through derivative weighting. This local perspective reveals regional heterogeneity that global methods obscure. Moreover, the augmented co-monotonicity embedding incorporating inter-feature associations enables discovery of feature modules with coherent relationships, information that pairwise correlations between outcome and individual features cannot capture.
7.4 Limitations and Future Directions
@@
Several limitations suggest directions for future development. The k-nearest neighbor graph construction introduces discrete approximation of the underlying manifold, with performance sensitive to the neighborhood size parameter $k$. While we employ principled selection based on connectivity and local geometry, the optimal $k$ remains data-dependent. Adaptive graph construction methods that locally vary neighborhood size or employ data-dependent distance metrics may improve approximation quality.
The current framework applies to continuous or binary outcomes through spectral filtering on the graph Laplacian. Extension to general discrete outcomes (count data, categorical responses) and survival outcomes (time-to-event data) requires development of appropriate smoothing operators that respect the outcome type while preserving the geometric structure. Preliminary work on Poisson regression suggests that weighted Laplacians with outcome-dependent masses can accommodate count responses, but comprehensive treatment awaits future investigation.
@@
Computational scaling presents challenges for very high-dimensional feature spaces. Computing the augmented co-monotonicity embedding requires $O(m^{2})$ coefficients for $m$ features, and the induced graph $G_{k}(\text{cm})$ in dimension $m(m+1)/2$ may suffer from curse of dimensionality when $m$ is large. Dimensionality reduction through principal component analysis on association profiles, or projection pursuit methods that identify discriminating directions in association space, may alleviate this burden. Alternatively, sparse formulations that compute only selected inter-feature associations could reduce dimensionality while preserving essential structure.
The path monotonicity validation criterion for long edges employs heuristic thresholds (quantile $q\in[0.75,0.90]$ for identifying long edges, threshold $\theta\in[0.85,0.95]$ for accepting monotonicity). While these ranges perform well empirically across diverse datasets in our experience, principled data-driven selection of these parameters through cross-validation or information criteria would strengthen the methodology. Theoretical analysis characterizing the relationship between threshold choice and basin structure properties could guide selection.
@@
Several promising extensions warrant exploration. The framework naturally accommodates multi-modal data integration: constructing co-monotonicity embeddings that incorporate associations within and between multiple feature sets (genomic, proteomic, metabolomic) enables joint analysis that leverages complementary information. Temporal extensions for longitudinal data could track how co-monotonicity cells evolve over time, identifying feature modules whose relationships with outcomes change with disease progression or treatment. Causal inference extensions employing directed graphs and structural equation models could move beyond association to intervention prediction, answering questions about which features to manipulate to alter outcomes.
7.5 Broader Impact
The geometric framework addresses methodological challenges that appear across scientific domains wherever high-dimensional structured data and context-dependent relationships arise. In microbiome research, the methods enable discovery of microbial community structures that associate with health outcomes, revealing that the same bacterial taxa may promote or protect against disease depending on the broader community context. In single-cell genomics, gradient flow cells could identify cell state trajectories while co-monotonicity cells discover gene modules with coordinated expression programs that vary across cell types. In spatial transcriptomics, the framework can partition tissue regions based on molecular signatures and identify context-dependent gene-phenotype relationships across anatomical structures. In ecological data, co-monotonicity analysis could reveal how species interactions vary across environmental gradients, with different competitive or mutualistic networks operating in different habitats.
Beyond specific applications, the framework demonstrates the value of geometry-first approaches to statistical inference on complex data. Rather than treating samples as unstructured collections of observations, we represent them through graphs that encode proximity relationships, equip these graphs with Riemannian structure that respects local density and geometry, and discover structure through gradient flow and association profiles that respect this geometry. This perspective bridges differential geometry, spectral graph theory, and statistical inference, providing a mathematically principled foundation for methods that adapt to intrinsic data structure rather than imposing ambient Euclidean geometry.
The Bayesian inference framework we develop addresses a conceptual limitation of classical hypothesis testing that becomes acute in exploratory high-dimensional analysis. Traditional $p$-values answer the question ”how surprising would the data be if no association existed?” but investigators typically care more about ”how strong is the evidence for meaningful association?” and ”how uncertain are we about effect magnitudes?” Posterior probabilities and credible intervals directly address these questions, enabling principled uncertainty quantification that propagates through the entire analysis pipeline from conditional expectation estimation through biclustering to cell-level coherence assessment. This approach aligns with growing recognition in statistics that effect sizes and uncertainty bounds provide more actionable information than binary significance decisions.
7.6 Conclusion
We have developed a comprehensive geometric framework for discovering and quantifying spatially heterogeneous associations in high-dimensional structured data. By representing samples through Riemannian graphs, employing path-validated gradient flow to partition based on outcome geometry, defining co-monotonicity coefficients that respect graph structure, and constructing association space embeddings that enable biclustering, the framework addresses limitations of global methods that assume homogeneous relationships and local methods that lack power to detect structure. The Bayesian posterior sampling approach provides rigorous uncertainty quantification that respects the geometry of the data manifold while enabling inference at multiple scales, from individual vertices to regional coherence measures.
The framework extends naturally to multi-modal data integration, temporal dynamics, and potentially causal inference, suggesting a broad research program for geometric statistical inference. While substantial work remains in implementation optimization, theoretical convergence analysis, and empirical validation across diverse application domains, the core methodological contributions presented here provide a foundation for this program. Detailed computational documentation, formal theoretical results, and comprehensive real-data case studies will be provided in subsequent versions of this manuscript.
The methods are implemented in the R package gflow, available at https://github.com/pgajer/gflow, enabling immediate application by researchers facing similar challenges in their own high-dimensional biomarker studies. We anticipate that the geometric perspective on association analysis will prove valuable across scientific domains where understanding context-dependent relationships is essential for mechanistic insight and intervention design.

### v2

@@ 4.6 Matrix Extension for Multivariate Analysis
In applications, we augment $\text{CM}(y,Z)$ with selected columns from $\text{CM}(Z,Z)$ to form an extended association profile. For a vertex $v$, this profile includes both $\text{cm}(y,z_{j})(v)$ for all features $j$ and $\text{cm}(z_{j},z_{k})(v)$ for pairs $(j,k)$ of interest. These profiles embed vertices into an association space where proximity reflects similarity in how outcome and features relate locally. Vertices with similar profiles form co-monotonicity cells, which we will explore in subsequent sections as a basis for geometric multiple testing and biclustering.
@@ 7. Discussion
We have developed a geometric decomposition framework for statistical inference on high-dimensional structured data, addressing heterogeneous associations through two complementary strategies: gradient flow cells based on outcome geometry, and co-monotonicity cells based on association profiles. The path monotonicity validation criterion resolves long edge artifacts in discrete gradient flow through geometric verification rather than heuristic thresholding. The co-monotonicity coefficients provide spatially-resolved alternatives to global correlation measures, with derivative weighting connecting to angular correlation of gradients in the continuous manifold limit.
Relationship to Existing Methods
@@
Our framework extends Morse-Smale regression [Gerber et al., 2012] through several key innovations: robust conditional expectation estimation on density-aware Riemannian graphs, systematic spurious extrema removal via prominence filtering and basin overlap clustering, path monotonicity validation for gradient flow edges, and co-monotonicity measures enabling discovery of association-driven structure complementary to outcome-driven partitions.
@@
The framework situates within geometric data analysis rather than topological data analysis. While TDA emphasizes topological invariants like persistent homology that remain unchanged under continuous deformations, our approach focuses on geometric properties like distances, angles, and geodesics dependent on specific metric structure. The Riemannian graphs with vertex and edge masses capture local density and geometry that spectral methods exploit. This geometric emphasis connects our work to manifold learning and spectral graph theory more than to persistent homology.
Limitations and Future Directions
@@
The k-nearest neighbor graph construction introduces discrete approximation of the underlying manifold, with performance sensitive to neighborhood size $k$. While we employ principled selection based on connectivity and local geometry, adaptive methods that locally vary neighborhood size may improve approximation quality.
@@
Computational scaling presents challenges for very high-dimensional feature spaces. Computing the augmented co-monotonicity embedding requires $O(m^{2})$ coefficients for $m$ features, potentially suffering from curse of dimensionality when $m$ is large. Dimensionality reduction through principal component analysis on association profiles or sparse formulations computing only selected inter-feature associations may alleviate this burden.
The current framework applies to continuous or binary outcomes through spectral filtering. Extension to general discrete outcomes (count data, categorical responses) and survival outcomes requires development of appropriate smoothing operators that respect the outcome type while preserving geometric structure.
@@
The framework naturally accommodates multi-modal data integration through co-monotonicity embeddings incorporating associations within and between multiple feature sets. Temporal extensions for longitudinal data could track how co-monotonicity cells evolve over time, while causal inference extensions employing directed graphs could move beyond association to intervention prediction.
Appendix: Hodge Laplacian Matrix Formula
Derivation of the Hodge Laplacian on chains formula.
The chain Hodge Laplacian on $p$-chains is defined as
$L_{p}=\partial_{p+1}\partial_{p+1}^{*}+\partial_{p}^{*}\partial_{p}$
Thus
$L_{p}:C_{p}\to C_{p}$
with
$\partial_{p+1}:C_{p+1}\to C_{p}$
$\partial_{p+1}^{*}:C_{p}\to C_{p+1}$
When inner products are represented by matrices $M_{p}$ and $M_{p-1}$, the adjoint takes the matrix form:
$\partial_{p}^{*}=M_{p}^{-1}\partial_{p}^{T}M_{p-1}$
To derive this, let $\alpha=\sum_{i}\alpha_{i}\sigma_{i}^{(p)}$ and $\beta=\sum_{j}\beta_{j}\sigma_{j}^{(p-1)}$ where $\{\sigma_{i}^{(p)}\}$ are basis $p$-simplices. In matrix notation:
$\langle\partial_{p}\alpha,\beta\rangle_{p-1}=(\partial_{p}\boldsymbol{\alpha})^{T}M_{p-1}\boldsymbol{\beta}=\boldsymbol{\alpha}^{T}\partial_{p}^{T}M_{p-1}\boldsymbol{\beta}$
$\langle\alpha,\partial_{p}^{*}\beta\rangle_{p}=\boldsymbol{\alpha}^{T}M_{p}(\partial_{p}^{*}\boldsymbol{\beta})$
Equating these for all $\boldsymbol{\alpha}$ gives $M_{p}(\partial_{p}^{*}\boldsymbol{\beta})=\partial_{p}^{T}M_{p-1}\boldsymbol{\beta}$, hence the matrix formula
$\partial_{p}^{*}=M_{p}^{-1}\partial_{p}^{T}M_{p-1}$
and hence
$\partial_{p+1}^{*}=M_{p+1}^{-1}\partial_{p+1}^{T}M_{p}$
Therefore
$\partial_{p+1}\partial_{p+1}^{*}=\partial_{p+1}M_{p+1}^{-1}\partial_{p+1}^{T}M_{p}$
and
$\partial_{p}^{*}\partial_{p}=M_{p}^{-1}\partial_{p}^{T}M_{p-1}\partial_{p}$
If we use the matrix notation $B_{p}$ for $\partial_{p}$, then
$\partial_{p+1}\partial_{p+1}^{*}=B_{p+1}M_{p+1}^{-1}B_{p+1}^{T}M_{p}$
and
$\partial_{p}^{*}\partial_{p}=M_{p}^{-1}B_{p}^{T}M_{p-1}B_{p}$